### v1

# OATS: Outlier-Aware Pruning through Sparse and Low Rank Decomposition

Abstract
The recent paradigm shift to large-scale foundation models has brought about a new era for deep learning that, while has found great success in practice, has also been plagued by prohibitively expensive costs in terms of high memory consumption and compute. To mitigate these issues, there has been a concerted effort in post-hoc neural network pruning techniques that do not require costly retraining. Despite the considerable progress being made, existing methods often exhibit a steady drop in model performance as the compression increases. In this paper, we present a novel approach to compressing large transformers, coined OATS, that utilizes the second moment information in the input embeddings to decompose the model weights into a sum of sparse and low-rank matrices. Without any retraining, OATS achieves state-of-the-art performance when compressing models by up to $60\%$ on large language models such as Llama-3 and Phi-3 and vision transformers such as ViT and DINOv2 while delivering up to 1.37×1.37$\times$1.37 × the CPU acceleration versus a model that was comparably pruned.
1 Introduction
Large scale transformer-based models have found great success in a number of domains ranging from image classification (Wu et al., 2020), language modelling (Devlin et al., 2019), and question answering (Brown et al., 2020). However, these models contain billions of parameters making them computationally expensive to train and deploy which has lead to an increased demand for resource-saving techniques like model quantization (Dettmers et al., 2022; Egiazarian et al., 2024), parameter efficient fine-tuning (Hu et al., 2022; Zhao et al., 2024b), and, most relevant to this work, neural network pruning (Frantar & Alistarh, 2023).
Pruning has been a key focus for model compression since the early days of deep networks (Mozer & Smolensky, 1988; LeCun et al., 1989; Hassibi & Stork, 1992). Over the years, various pruning techniques have emerged, introducing sparsity either before training (Lee et al., 2019; Wang et al., 2020; Tanaka et al., 2020; de Jorge et al., 2021), during training (Zhu & Gupta, 2018; Evci et al., 2020), or post-training (Benbaki et al., 2023). In the context of large foundation models, post-training pruning methods, particularly those requiring minimal (Xia et al., 2022; Ma et al., 2023) or no re-training (Frantar & Alistarh, 2023; Sun et al., 2024b; Ashkboos et al., 2024; Zhang et al., 2024b) are preferred for their computational efficiency. These techniques, when compressing models by $50\%$, have demonstrated the ability to accelerate end-to-end CPU inference by up to 1.8×1.8$\times$1.8 × (Frantar & Alistarh, 2023; Yin et al., 2024b) and GPU inference by up to 1.63×1.63$\times$1.63 × using structured N:M sparsity (Mishra et al., 2021), highlighting their potential in real-world applications.
Despite the significant advancements in pruning techniques, it was recently shown that current methods suffer from a consistent degradation in model performance as compression levels increase (Yin et al., 2024a). Moreover, although structured pruning offers greater potential for acceleration compared to unstructured pruning, it often imposes a much steeper trade-off in terms of model accuracy and effectiveness (Chen et al., 2022). These challenges underscore the need for more sophisticated pruning strategies to strike a better balance between efficiency and performance.
1.1 Contributions
To mitigate these issues, we introduce Outlier-Aware Pruning Through Sparse and Low Rank Decomposition (OATS), a novel retraining-free method for compressing large transformers that utilizes the second moment of the input embeddings to approximate the model’s weight matrices as a sum of a sparse matrix and a low-rank matrix.
We evaluate OATS on recent large language models – Phi-3 (Abdin et al., 2024) and Llama-3 (Dubey et al., 2024) – and vision transformers – ViT (Wu et al., 2020) and DinoV2 (Oquab et al., 2023) – demonstrating that OATS achieves new state-of-the-art performance across a wide range of commonly employed performance metrics. Furthermore, by combining structured pruning with unstructured pruning, our benchmarks utilizing the Deepsparse Engine by NeuralMagic (2021) reveal that OATS leads to more CPU speed-up across all levels of compression when compared to models that utilize just unstructured pruning.
To facilitate the interpretability of the sparse and low-rank terms found by OATS, we run experiments where we divide the compressed model into two models, a sparse model and a low-rank model, and visualize their respective attention heat maps utilizing attention rollout (Abnar & Zuidema, 2020). Our visualizations on vision transformers (Wu et al., 2020) reveal a complementary relationship between the two models, with each focusing on different key areas of the image, effectively segmenting it into distinct regions.
2 The OATS Algorithm
The key observation behind the OATS algorithm is that the weight matrices, ${\bm{W}}\in\mathbb{R}^{d_{out}\times d_{in}}$, in a transformer model can be faithfully approximated as a summation of a sparse and low rank matrix by solving the following optimization problem, commonly known as Robust PCA (Chandrasekaran et al., 2011; Candès et al., 2011):
$\min_{{\bm{S}},{\bm{L}}\in\mathbb{R}^{d_{out}\times d_{in}}}\|{\bm{W}}-{\bm{S}}-{\bm{L}}\|_{F}^{2}\;\text{ s.t. }\;\text{Rank}({\bm{L}})\leq r,\;\|{\bm{S}}\|_{0}\leq k.$ (1)
2.1 Alternating Thresholding
To solve Equation 1, OATS leverages the alternating thresholding algorithms proposed by Netrapalli et al. (2014) and Bertsimas et al. (2024) that iteratively alternates between solving for the low-rank term ${\bm{L}}$, through singular-value thresholding, and for the sparse term ${\bm{S}}$, through hard-thresholding. Given a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, singular-value thresholding, also known as truncated SVD, is defined as:
$\textsc{TruncatedSVD}({\bm{A}},r)={\bm{U}}_{r}\mathbf{\Sigma}_{r}{\bm{V}}_{r}^{\top},$
where ${\bm{U}}_{r},\mathbf{\Sigma}_{r},{\bm{V}}_{r}^{\top}$ correspond to the matrices formed by retaining only the top-$r$ singular vectors and singular values from the full SVD of ${\bm{A}}$. Hard-thresholding, which succeeds the singular-value thresholding step, is defined as:
$\textsc{HardThreshold}({\bm{A}},k)={\bm{M}}\odot{\bm{A}},$
Algorithm 1 AlternatingThreshold
1:  Inputs:
2:      Weight Matrix: ${\bm{W}}\in\mathbb{R}^{d_{out}\times d_{in}}$
3:      Iterations: $N$
4:      Rank: $r$
5:      Nonzeros: $k$
6:  Procedure:
7:  ${\bm{S}}=\mathbf{0}$
8:  for $t=1$ to $N$ do
9:     ${\bm{L}}=\textsc{TruncatedSVD}({\bm{W}}-{\bm{S}},r)$
10:     ${\bm{S}}=\textsc{HardThreshold}({\bm{W}}-{\bm{L}},k)$
11:  end for
12:  return: ${\bm{S}}$, ${\bm{L}}$
where ${\bm{M}}\in\mathbb{R}^{m\times n}$ is a binary matrix with $k$ non-zero entries coinciding with the $k$ largest entries in magnitude in ${\bm{A}}$. Alternatively, the hard-thresholding can also be performed row-wise rather than layer-wise in which case ${\bm{M}}$ would be a binary matrix with $m\cdot\left\lfloor\frac{k}{m}\right\rfloor$ non-zero entries coinciding with the $\left\lfloor\frac{k}{m}\right\rfloor$ largest entries in magnitude in each row of ${\bm{A}}$.
These steps are summarized in Algorithm 1 on the right. To optimize memory usage, the low-rank term ${\bm{L}}$ is stored as its two low-rank components: ${\bm{U}}_{r}$ and $\mathbf{\Sigma}_{r}{\bm{V}}_{r}^{\top}$.
2.2 Incorporating Outlier Information
The alternating thresholding on its own yields suboptimal results because the activations of large-scale transformers exhibit a small number of large-magnitude features and altering these negatively impacts model performance (Kovaleva et al., 2021; Dettmers et al., 2022; Sun et al., 2024a). OATS takes inspiration from Wanda (Sun et al., 2024b) and incorporates outlier information through a diagonal scaling matrix ${\bm{D}}\in\mathbb{R}^{d_{in}\times d_{in}}$ that captures the second moment of the input activations
${\bm{D}}=\text{diag}\left(\sqrt{{\bm{X}}^{\top}{\bm{X}}}\right),$
where ${\bm{X}}\in\mathbb{R}^{B\times d_{in}}$ and $B$ is the product of the batch size and sequence length. The weight matrix ${\bm{W}}$ is then scaled by ${\bm{D}}$ before applying AlternatingThreshold:
${\bm{S}},{\bm{L}}=\textsc{AlternatingThreshold}\left({\bm{W}}{\bm{D}},N,r,k\right).$
To undo the scaling, OATS applies the inverse transformation to reach the compressed weight
${\bm{W}}_{\text{compressed}}=({\bm{S}}+{\bm{L}}){\bm{D}}^{-1},$
where it leverages the fact that ${\bm{D}}$ is diagonal so that it both preserves the sparsity pattern of ${\bm{S}}$ and is easy to invert. Aligned with (Frantar & Alistarh, 2023; Sun et al., 2024b), and Zhang et al. (2024b), the activations are calculated through a calibration set that is propagated through the compressed layers.
2.3 OATS Parameters
To determine the rank $r$ and the number of nonzeros $k$, OATS takes in as input two hyperparameters: the compression rate, $\rho\in(0,1)$, and the rank ratio, $\kappa\in(0,1)$. The compression rate coincides with the sparsity rate required by existing pruning algorithms and is defined as:
$\rho=1-\frac{\#\text{ of nonzero parameters in compressed layer}}{\#\text{ of parameters in original layer}}=1-\frac{k+r(d_{out}+d_{in})}{d_{out}\cdot d_{in}}.$
The rank ratio represents the proportion of nonzero parameters that appear in the low-rank term:
$\kappa=\frac{r(d_{out}+d_{in})}{(1-\rho)d_{out}\cdot d_{in}}.$
Given a fixed compression rate $\rho$ and rank ratio $\kappa$, the two equations above can be solved to obtain the rank $r$ and nonzeros $k$:
$r=\left\lfloor\kappa\cdot(1-\rho)\cdot\frac{d_{out}\cdot d_{in}}{d_{out}+d_{in}}\right\rfloor\qquad k=\left\lfloor(1-\kappa)\cdot(1-\rho)\cdot d_{out}\cdot d_{in}\right\rfloor.$ (2)
The complete OATS algorithm pseudocode can be found in Algorithm 2 below.
Algorithm 2 OATS
1:  Inputs:
2:      Layer Inputs: ${\bm{X}}^{\ell}\in\mathbb{R}^{B\times d_{in}}$
3:      Layer Matrix: ${\bm{W}}^{\ell}\in\mathbb{R}^{d_{out}\times d_{in}}$
4:      Compression Rate: $\rho$
5:      Rank Ratio: $\kappa$
6:      Iterations: $N$
7:  Procedure:
8:  $r\leftarrow\left\lfloor\kappa\cdot(1-\rho)\cdot\frac{d_{out}\cdot d_{in}}{d_{out}+d_{in}}\right\rfloor,\;k\leftarrow\left\lfloor(1-\kappa)\cdot(1-\rho)\cdot d_{out}\cdot d_{in}\right\rfloor$
9:  ${\bm{D}}\leftarrow\text{diag}(\sqrt{{\bm{X}}^{\top}{\bm{X}}})$
10:  ${\bm{L}},{\bm{S}}\leftarrow\textsc{AlternatingThreshold}({\bm{W}}{\bm{D}},N,r,k)$
11:  ${\bm{W}}\leftarrow({\bm{L}}+{\bm{S}}){\bm{D}}^{-1}$
12:  return: ${\bm{X}}^{\ell+1}\leftarrow{\bm{W}}{\bm{X}}^{\ell}$
3 Experiments on Large Language Models
3.1 Experiment Setup
Models and Tasks
We evaluate OATS on two state-of-the-art families of large language models: Phi-3 (Abdin et al., 2024) and Llama-3 (Dubey et al., 2024). To gauge the algorithm’s performance under various model sizes, we select Phi-3 Mini, a $3.8$B parameter model, Phi-3 Medium, a $14$B parameter model, and Llama-3 8B, an $8$B parameter model, so that models of various sizes are represented in our experiments. We utilize LM Harness developed by Gao et al. (2024) to evaluate zero-shot performance on eight tasks, five-shot performance on the Massive Multitask Language Understanding benchmark by Hendrycks et al. (2021), and language generation on WikiText-2.
Parameters Phi-3 Llama-3
Iterations 80 80
Rank Ratio 25% 30%
Table 1: Hyperparameters utilized for OATS across model families. Both parameters are further ablated in Section 3.3.
Pruning Benchmarks
As OATS does not require costly retraining after model compression, we opt to benchmark it with three current state-of-the-art algorithms that similarly do not require such overhead: SparseGPT by Frantar & Alistarh (2023), Wanda by Sun et al. (2024b), and DSNoT [^1] DSNoT experiments are run with both SparseGPT and Wanda. We report the best results across the two. Further details are in Appendix A.4. by Zhang et al. (2024b). The parameters utilized for OATS are depicted in Table 1.
Calibration Data
Remaining consistent with Frantar & Alistarh (2023), Sun et al. (2024b), and Zhang et al. (2024b), our calibration data consists of 128 sequences of length 2048 sampled from the first shard of the C4 training set (Raffel et al., 2020). To ensure consistency, we utilize the same calibration data for all pruning algorithms that we benchmark.
Layer-Wise Compression Rates
We benchmark our algorithm across a wide range of compression rates: {0.3, 0.4, 0.5, 0.6}. For compression rates at or below $0.5$, we compress all transformer blocks uniformly. At the higher compression rate of $0.6$, we utilize Outlier Weighed Layerwise Sparsity Ratios (OWL) developed by Yin et al. (2024b) which was shown to lead to significant performance improvements at higher compression rates. We exclude pruning any linear layers that are present in the model head and embeddings which conforms with what was done in the prior works by Frantar & Alistarh (2023), Sun et al. (2024b), and Zhang et al. (2024b).
Hardware Speedup
To measure the CPU speedup generated by the unstructured sparsity and low-rank compression of OATS, we utilize the DeepSparse Inference Engine developed by NeuralMagic (2021). More recently, NVIDIA’s sparse tensor cores are able to leverage for acceleration structured N:M sparsity, a sparsity pattern where every $N$ in $M$ parameters are zero (Mishra et al., 2021). We include in our experiments the natural extension of OATS to structured N:M sparsity where we impose the structured sparsity pattern onto the sparse term in the decomposition.
3.2 Results
Five-shot MMLU
Depicted in Table 2 below is the MMLU accuracy of OATS relative to current state of the art pruning algorithms. OATS is able to outperform all prior methods across all compression rates with an increasing gap as the compression rate increases. Specifically, at $50\%$ compression, OATS surpasses previous pruning algorithms by a margin of $2.52\%$ on Phi-3 Medium and $5.42\%$ on Phi-3 Mini.
Compression Rate Method Phi-3 Llama-3 8B
Mini Medium
0% Dense 70.34 76.78 64.97
30% SparseGPT 68.31 74.12 64.25
Wanda 67.63 75.18 63.67
DSNoT 68.02 75.13 63.72
OATS 68.84 76.15 65.22
40% SparseGPT 63.47 72.42 60.91
Wanda 64.15 73.34 60.33
DSNoT 63.57 73.20 59.99
OATS 65.75 74.99 62.46
50% SparseGPT 53.22 67.63 53.60
Wanda 54.57 69.76 49.83
DSNoT 54.28 68.65 49.20
OATS 59.99 72.28 56.46
Table 2: Comparison of average five-shot accuracies (%) on MMLU under different compression rates.
Zero-shot Tasks
Table 3 below reports the zero-shot accuracy of OATS relative to current state of the art pruning algorithms averaged across the following eight popular tasks: PIQA (Bisk et al., 2020); HellaSwag (Zellers et al., 2019); Winogrande (Sakaguchi et al., 2021); OpenBookQA (Mihaylov et al., 2018); RTE (Wang et al., 2018); BoolQ (Clark et al., 2019); ARC-e and ARC-c (Clark et al., 2018). Mirroring the trend observed in the five-shot results, the improvement from OATS compared to prior pruning algorithms increases with compression, culminating in a $2.05\%$ advantage over prior methods at $50\%$ compression for Phi-3 Mini.
Compression Rate Method Phi-3 Llama-3 8B
Mini Medium
0% Dense 71.99 74.27 69.79
30% SparseGPT 70.63 74.53 69.08
Wanda 70.66 74.05 68.63
DSNoT 71.20 74.03 68.98
OATS 71.48 74.04 69.34
40% SparseGPT 69.18 74.40 67.58
Wanda 68.80 73.01 67.04
DSNoT 69.08 72.90 66.65
OATS 70.04 74.46 68.68
50% SparseGPT 66.36 73.25 64.66
Wanda 65.03 70.96 63.27
DSNoT 65.33 71.12 62.74
OATS 68.41 73.39 65.71
Table 3: Comparison of average zero-shot accuracies (%) under different compression rates. Task-specific scores can be found in Appendix A.1.
Generation Task
Depicted in Table 4 below is the WikiText-2 perplexity of OATS relative to current state of the art pruning algorithms. At $50\%$ compression, OATS results in an $8.5\%$ reduction in perplexity on the larger Phi-3 Medium model, and an even larger $9\%$ reduction on Phi-3 Mini and Llama-3 8B.
Compression Rate Method Phi-3 Llama-3 8B
Mini Medium
0% Dense 9.50 6.21 10.17
30% SparseGPT 11.19 7.48 9.71
Wanda 10.71 7.28 9.39
DSNoT 10.51 7.11 9.36
OATS 10.27 6.85 9.59
40% SparseGPT 13.03 8.52 10.01
Wanda 12.59 8.49 9.74
DSNoT 12.17 8.24 9.60
OATS 11.53 7.70 9.24
50% SparseGPT 16.80 9.89 11.95
Wanda 17.23 10.12 12.36
DSNoT 16.68 9.96 12.41
OATS 15.18 9.05 10.87
Table 4: Comparison of perplexity (lower is better) on WikiText-2 under different compression rates.
Method Phi-3 Mini Llama-3 8B
SparseGPT 46.20 39.48
Wanda 44.22 31.20
DSNoT 44.75 33.28
OATS 52.59 44.46
[…]
Performance Under High Compression
[…]
3.3 Ablations for OATS
We conduct ablation studies for OATS on Phi-3 Mini at $40\%$ compression rate with a rank ratio of $20\%$ to depict the quantitative impact of the following design choices that were made:
  • Scaling the weights by the second moment of the input activations, ${\bm{D}}$, versus not scaling.
  • Pruning the weights per each output row in the matrix versus pruning layer-wise.
The results of our ablations are depicted in Table 6 below:
Ablation MMLU ()($\uparrow$)( ↑ ) Zero-shot ()($\uparrow$)( ↑ ) Perplexity ()($\downarrow$)( ↓ )
No Scaling Layer-Wise 62.46 67.58 19.21
Row-Wise 65.31 68.22 18.34
Scaling by ${\bm{D}}$ Layer-Wise 64.44 70.52 11.68
Row-Wise 65.84 70.71 11.50
Table 6: Ablations of OATS on Phi-3-Mini at $40\%$ compression rate with a rank ratio of $20\%$. The experiments show that scaling the weights by the second moment of the input activations and pruning row-wise significantly improves the performance of OATS.
In addition to the ablations, in Figure 1 below, we show how the rank ratio and the number of iterations impact the performance of OATS .
0.10.20.30.40.50.60.74040404050505050606060607070707080808080Rank RatioAccuracy (%)120406080100120140160Number of Iterations0-shotMMLU
Figure 1: The effect of varying the rank ratio and number of iterations on zero-shot and five-shot accuracy.
The experiments reveal that the performance under the rank ratio generally saturates between $25\%$ to $30\%$ prior to deteriorating. For the number of iterations, there is a sharp increase in performance from 00-$20$ iterations before leveling off and saturating at around $80$ iterations.
3.4 Hardware Speedup
CPU Speedup
We utilize the DeepSparse engine by NeuralMagic (2021) to measure the CPU acceleration induced by OATS compared to pruned models with unstructured sparsity. We run end-to-end inference on a compressed Phi-3 Medium 15B model for a single batch of 2048 tokens on an Intel Xeon Gold 6148 CPU @ 2.40GHz with 32 cores. The achieved speed-ups are depicted in Table 7 below.
Compression Rate Method Throughput (tokens/sec) Speedup
Dense 4.03 1.00×1.$0$
30% Unstructured Pruning 4.32 1.07×1.07$\times$1.07 ×
OATS 5.58 $\mathbf{1.38\times}$
40% Unstructured Pruning 5.08 1.26×1.26$\times$1.26 ×
OATS 6.86 $\mathbf{1.73\times}$
50% Unstructured Pruning 7.16 1.78×1.78$\times$1.78 ×
OATS 8.31 $\mathbf{2.06\times}$
Table 7: Comparison of throughput and speedup achieved through OATS and unstructured pruning methods relative to their dense counterparts.
Our results highlight that by trading unstructured sparsity for structured sparsity through the low-rank terms, OATS leads to significantly more CPU speed-up compared to models that rely solely on unstructured pruning. In particular, at $40\%$ compressed, OATS has 1.37×1.37$\times$1.37 × the speed-up than unstructured pruning.
N:M Performance
We benchmark the N:M performance of OATS against the 2:4 structured sparsity performance of current state of the art pruning algorithms. Since the low-rank terms found by OATS remains dense, we opt to impose the sparser structured sparsity pattern of 2:8 on the sparse term returned by OATS testing various rank ratios of $0.25$, $0.3$, $0.35$, $0.4$, $0.45$, and $0.5$. In Figure 2 below, we depict the compression ratio versus the 5-shot MMLU accuracy of different compression algorithms.
5054586164663030303040404040MMLU Avg. AccuracyPhi-3 Mini505458616466Llama-3 8BOATSSparseGPTWandaDSNoT
Figure 2: Experiments evaluating OATS with 2:8 structured sparsity on the sparse terms against 2:4 sparsity of state-of-the pruning algorithms. The rank ratio for OATS is varied to capture the performance across different compression rates.
Our results show that despite having a much stricter structured sparsity pattern of 2:8, OATS is able to recover the model performance through the presence of its low-rank terms. Specifically, at a fixed compression rate of $50\%$, OATS is able to outperform all prior state-of-the-art by up to $9.2\%$. In the case of Llama-3 8B, OATS surpasses previous methods by $2.86\%$ with $8\%$ more compression.
4 Experiments on Vision Transformers
We run experiments on Google’s ViT-Base (Wu et al., 2020), an $86.6$M parameter model trained in a supervised manner on ImageNet-21k (Ridnik et al., 2021) and fine-tuned on ImageNet 2012 (Russakovsky et al., 2015), and DinoV2-Giant (Oquab et al., 2023), a $1.14$B parameter model that was trained utilizing self-supervised learning.
We benchmark OATS against the same three pruning algorithms: SparseGPT, Wanda, and DSNoT, by evaluating on the validation split of ImageNet (Russakovsky et al., 2015). A subset of size $2048$ from the training split of ImageNet is utilized for the calibration data and is maintained consistent across all pruning experiments. All OATS experiments utilize a rank ratio of $\kappa=20\%$ and $N=80$ iterations. We exclude from compression the embedding layers and the classifier layer.
Depicted in Table 8 below are the validation accuracies of ImageNet for ViT-Base and DinoV2-Giant under various compression rates.
Compression Rate Method ViT-Base DinoV2-Giant
$0\%$ Dense 80.33 86.55
30% SparseGPT 80.21 86.46
Wanda 80.28 86.47
DSNoT 80.16 86.46
OATS 80.15 86.52
40% SparseGPT 79.58 86.39
Wanda 79.34 86.32
DSNoT 79.46 86.37
OATS 79.86 86.46
50% SparseGPT 78.44 86.04
Wanda 76.19 85.81
DSNoT 76.90 85.93
OATS 78.77 86.14
[…]
5 Visualizing and Interpreting the Decomposition
To develop a better understanding of how the sparse and low rank components individually contribute to the flow of information through the attention mechanism, we compute and visualize the attention rollout (Abnar & Zuidema, 2020) of the compressed model when:
  • All low-rank terms are set to zero and inputs are propagated through only the sparse terms.
  • All sparse terms are set to zero and inputs are propagated through only the low-rank terms.
Figure 3 below provides a visualization of how the information would flow through a standard transformer block for both settings.
[…]
(a) Sparse Terms Only
[…]
(b) Low-Rank Terms Only
Figure 3: A visualization of how the attention rollout is computed to isolate the contribution of the sparse terms versus low-rank terms given by the OATS compression algorithm.
We utilize various images in the Microsoft COCO dataset (Lin et al., 2014) to generate the attention rollout. Figure 4 below depicts the attention rollout computed using the scheme described above on ViT-B that was compressed by $50\%$ with a rank ratio of $20\%$.
Original: Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Sparse: Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-Rank: Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Attention rollout visualization applied to various images on the Microsoft COCO dataset.
The rollout visualizations show that the sparse and low-rank terms capture distinct areas of the image, effectively segmenting it, and reveal three distinct patterns in how the image is partitioned. The first type of behavior, which is also commonly exhibited in the classical setting (Candès et al., 2011), is when one component (commonly sparse) captures the subject(s), while the other component (commonly low-rank) captures the background. The second behavior occurs when both components focus on different parts of the same subject, capturing distinct features. The third behavior arises when the image contains multiple subjects, with each component isolating a different subject. While these patterns provide initial insights into how the components process visual information, further investigation is needed to fully understand the underlying mechanisms driving these behaviors.
6 Related Works
Connection with Wanda
OATS utilizes the same outlier scaling as the the one that is employed by Wanda (Sun et al., 2024b). In fact, Wanda can be seen as a special case of OATS when the rank ratio $\kappa=0$, which, according to Equation 2, results in the low-rank term becoming the $\mathbf{0}$ matrix. This leads to OATS performing a singular hard thresholding step that is equivalent to the pruning step described by Wanda and returning a compressed matrix equal to:
${\bm{W}}_{\text{compressed}}=\textsc{HardThreshold}({\bm{W}}{\bm{D}},k){\bm{D}}^{-1}.$
Sparse and Low-Rank Approximation in Transformers
The emergence of sparse and low-rank structures in transformers has recently become an area of both theoretical and practical interest. On the theoretical front, Zhao et al. (2024c) showed that the logits of LLMs trained utilizing next token prediction converge to a low rank and sparse structure. On the practical front, Scatterbrain proposed by Chen et al. (2021) shows that its possible to approximate the entire attention mechanism with a single sparse and low rank representation. LoRAP by Li et al. (2024) finds a low-rank representation for the attention matrices by utilizing a similar scaling technique as OATS. Recent works, such as LoSparse (Li et al., 2023), LoRAPrune (Zhang et al., 2024a), and APT (Zhao et al., 2024a), propose variations of applying structured pruning on the weights while incorporating a low-rank adapter that is trained via gradient descent. We want to highlight that unlike OATS, these works utilize training and low-rank adaptation to recover model performance from structured pruning, whereas OATS depends on an intrinsic sparse and low-rank approximation of the weight matrices.
Robust PCA Algorithms
While OATS utilizes an alternating thresholding approach for its efficiency and simplicity, the search for fast Robust PCA algorithms has been a key area of interest since the inception of the problem. An alternative approach is to apply a convex relaxation, replacing the sparsity and low-rank constraints with an $\ell_{1}$ and nuclear norm surrogate leading to Stable Principle Component Pursuit (Zhou et al., 2010). Another way to enhance computational efficiency is to utilize gradient descent by parameterizing the low-rank matrix as ${\bm{L}}={\bm{U}}{\bm{V}}^{\top}$, and then performing gradient descent on ${\bm{U}}$ and ${\bm{V}}$ (Yi et al., 2016; Tong et al., 2021).
Pruning and Interpretability
While pruning has been proven to be a successful approach to model compression, there is still ongoing research into what pruning is pruning and how it impacts model performance. Paganini (2020) show that magnitude-based pruning has a disproportionate negative effect on underrepresented classes. In a similar vein, Yin et al. (2024a) show that pruning LLMs can lead to irreversible damage to model performance on tasks that are more challenging. We highlight that the low-rank term present in OATS might be able to mitigate the negative impacts of pruning raised by the prior works. Indeed, in Tables 2 and 5, we can see that at higher compression rates, the gap between OATS and prior state-of-the-art is significantly larger suggesting that the low-rank term plays a critical role in mitigating the loss in performance at high compression.
7 Conclusion
We have introduced OATS, an algorithm that distinguishes itself from prior algorithms by utilizing outlier information to determine a compressed sparse and low-rank decomposition without any re-training. We run a number of experiments showing that OATS is able to consistently outperform prior state-of-the-art across multiple benchmarks and compression rates while also improving on speed-up. Furthermore, in the search for interpretability, we have generated visualizations highlighting that the sparse and low-rank terms capture different segments of the image. We hope that our work can be utilized to investigate how sparse and low-rank structures in transformers can be fully leveraged for maximal memory and compute efficiency.
Appendix A Appendix
A.1 Zero-Shot Task-Specific Performance
Table 9 depicts the task-specific performance for the zero-shot evaluation results presented in Section 3.2 and Appendix A.4.
Model Compression Rate Method PIQA HellaSwag WinoGrande OpenBookQA RTE BoolQ ARC-e ARC-c Phi-3 Mini 0% Dense 81.23 77.50 73.56 46.80 75.81 85.32 78.45 57.25 30% SparseGPT 78.94 76.94 69.85 49.60 73.29 84.13 76.39 55.89 Wanda 79.65 76.27 71.59 48.00 73.65 83.70 77.23 55.20 DSNoT w/ SparseGPT 80.09 75.61 72.22 47.40 74.37 84.22 77.86 54.69 DSNoT w/ Wanda 80.41 75.52 72.06 47.60 74.37 84.53 79.55 55.55 OATS 80.03 77.07 72.61 47.60 74.37 84.92 77.44 57.76 40% SparseGPT 78.35 75.07 68.59 47.00 72.20 83.67 75.29 53.24 Wanda 78.35 73.87 69.30 45.40 71.84 83.18 76.52 51.96 DSNoT w/ SparseGPT 78.56 72.99 70.64 46.20 70.40 82.72 76.52 52.82 DSNoT w/ Wanda 79.33 73.06 70.88 44.00 70.76 83.79 77.40 53.41 OATS 79.38 75.86 70.01 46.60 72.56 83.98 76.85 55.12 50% SparseGPT 77.20 70.63 66.46 45.20 70.76 83.06 70.58 47.01 Wanda 76.33 67.70 66.38 41.80 66.43 81.83 72.43 47.35 DSNoT w/ SparseGPT 76.28 67.16 65.90 42.20 63.90 81.56 72.90 48.04 DSNoT w/ Wanda 75.52 66.54 67.64 43.00 65.34 82.54 73.48 48.55 OATS 77.26 71.64 69.53 44.80 73.65 81.28 77.10 52.05 Phi-3 Medium 0% Dense 81.66 82.83 75.85 50.00 77.62 88.17 78.41 59.64 30% SparseGPT 81.39 82.02 75.77 50.80 77.26 87.80 80.05 61.18 Wanda 81.39 80.88 76.01 49.40 76.90 87.74 79.59 60.49 DSNoT w/ SparseGPT 81.94 80.76 76.95 48.40 75.81 87.65 79.25 59.81 DSNoT w/ Wanda 81.66 81.03 77.27 49.20 76.53 87.80 78.96 59.81 OATS 81.07 82.09 74.43 51.20 78.34 88.38 78.16 58.70 40% SparseGPT 80.41 80.70 75.53 51.20 77.26 88.32 81.23 60.58 Wanda 79.87 78.15 75.45 48.60 77.26 87.71 78.11 58.96 DSNoT w/ SparseGPT 79.82 78.07 75.37 47.00 76.53 87.98 77.31 58.19 DSNoT w/ Wanda 80.30 78.11 74.66 47.80 77.26 88.04 78.11 58.87 OATS 81.39 81.72 75.06 51.00 77.62 87.65 80.39 60.84 50% SparseGPT 79.71 78.27 73.64 50.40 75.45 87.09 82.03 59.39 Wanda 78.29 74.07 74.03 45.00 75.81 85.72 77.44 57.34 DSNoT w/ SparseGPT 79.27 74.30 74.59 44.40 76.90 85.26 77.69 56.57 DSNoT w/ Wanda 78.56 73.81 75.14 43.60 75.81 86.33 77.53 58.02 OATS 81.07 79.18 76.09 50.20 74.73 87.77 80.05 58.02 Llama-3 8B 0% Dense 80.74 79.16 73.40 45.0 0 67.87 80.98 77.69 53.50 30% SparseGPT 80.36 78.58 73.24 44.40 66.79 81.38 76.81 51.11 Wanda 79.98 78.00 73.64 44.40 64.26 81.62 76.18 50.94 DSNoT w/ SparseGPT 80.20 78.12 73.80 44.40 65.70 82.20 75.72 51.71 DSNoT w/ Wanda 79.82 77.99 73.09 44.80 63.18 81.80 77.06 51.37 OATS 80.03 78.75 73.64 45.20 66.06 81.13 76.94 52.99 40% SparseGPT 79.16 76.74 73.32 41.80 64.26 81.31 74.71 49.32 Wanda 78.73 75.90 72.22 44.40 63.18 80.46 72.31 49.15 DSNoT w/ SparseGPT 78.29 75.92 73.32 42.60 58.48 80.86 73.11 47.70 DSNoT w/ Wanda 78.51 75.52 73.24 43.80 61.73 80.70 72.01 47.70 OATS 79.71 77.18 74.19 43.80 67.51 82.39 74.92 49.74 50% SparseGPT 77.58 73.12 72.85 40.80 59.21 79.30 69.28 45.14 Wanda 77.53 69.34 70.24 40.00 61.73 76.57 66.96 43.77 DSNoT w/ SparseGPT 76.88 69.45 69.30 39.60 59.21 77.25 67.93 43.32 DSNoT w/ Wanda 77.09 68.57 69.77 38.60 57.76 76.27 67.34 43.43 OATS 77.75 73.17 71.74 41.00 64.98 79.66 72.35 45.05
Table 9: Task-Specific Zero-Shot Results
A.2 MMLU Subjects
We evaluate on the following MMLU subjects:
  • Abstract Algebra
  • Business Ethics
  • College Computer Science
  • College Mathematics
  • Conceptual Physics
  • Formal Logic
  • Machine Learning
  • Miscellaneous
  • Philosophy
  • Global Facts
which aligns with the subset utilized in the codebase of Ashkboos et al. (2024) that can be found here: https://github.com/microsoft/TransformerCompression.
A.3 Attention Rollout: Details
To generate the attention rollout visualizations depicted in Section 5, we average the attention matrices across the attention heads and discard the bottom $40\%$ attention pixels. The act of discarding the lowest value attention pixels was inspired by the following blog post by Gil (2021).
A.4 Pruning Implementation and Hyperparameters
Our code and pruning method implementations are based on the following codebases:
We utilize Huggingface’s Transformers library to implement the large language models and vision transformers for our experiments (Wolf et al., 2020).
A.4.1 SparseGPT Hyperparameters
We utilize a blocksize of 128 across all experiments and a Hessian dampening of $0.01$ and $0.1$ where the ladder is utilized only when faced with non-positive definiteness issues related with the Cholesky decomposition.
A.4.2 DSNoT Hyperparameters
We run experiments utilizing DSNoT where the initial masks are generated by SparseGPT and Wanda. All DSNoT experiments were run with $50$ iterations and an update threshold of $0.1$. Depicted in Table 10 below show the results distinguishing between the two initial methods that were utilized.
Model Compression Rate Method MMLU ($\uparrow$) Avg. Zero-Shot($\uparrow$) Perplexity($\downarrow$) Phi-3 Mini 30% DSNoT w/ SparseGPT 67.01 70.81 10.55 DSNoT w/ Wanda 68.02 71.20 10.51 40% DSNoT w/ SparseGPT 62.94 68.86 12.29 DSNoT w/ Wanda 63.57 69.08 12.17 50% DSNoT w/ SparseGPT 53.99 64.74 16.71 DSNoT w/ Wanda 54.28 65.33 16.68 Phi-3 Medium 30% DSNoT w/ SparseGPT 74.89 73.82 7.11 DSNoT w/ Wanda 75.13 74.03 7.11 40% DSNoT w/ SparseGPT 73.15 72.54 8.24 DSNoT w/ Wanda 73.20 72.90 8.27 50% DSNoT w/ SparseGPT 68.65 71.12 9.96 DSNoT w/ Wanda 68.12 71.10 10.02 Llama-3 8B 30% DSNoT w/ SparseGPT 62.99 68.98 9.37 DSNoT w/ Wanda 63.72 68.64 9.36 40% DSNoT w/ SparseGPT 58.97 66.28 9.60 DSNoT w/ Wanda 59.99 66.65 9.68 50% DSNoT w/ SparseGPT 49.15 62.74 12.41 DSNoT w/ Wanda 49.20 62.35 12.42
Table 10: LLM performance metrics of DSNoT with different initial methods.
Table 11 below depicts the analogous results but for our vision transformer experiments:
Model Compression Rate Method Accuracy (%) ViT-Base 30% DSNoT w/ SparseGPT 80.01 DSNoT w/ Wanda 80.16 40% DSNoT w/ SparseGPT 79.12 DSNoT w/ Wanda 79.46 50% DSNoT w/ SparseGPT 75.83 DSNoT w/ Wanda 76.90 DinoV2-Giant 30% DSNoT w/ SparseGPT 86.46 DSNoT w/ Wanda 86.45 40% DSNoT w/ SparseGPT 86.37 DSNoT w/ Wanda 86.30 50% DSNoT w/ SparseGPT 85.87 DSNoT w/ Wanda 85.93
Table 11: ImageNet Validation Accuracy of DSNoT with different initial methods.